# Channel Estimation for XL-MIMO Systems with Decentralized Baseband Processing: Integrating Local Reconstruction with Global Refinement

### v2

@@ IV-B Proposed SBL-GNNs Estimator
Utilizing the Lemma 1, the precision vector $\boldsymbol{\gamma}_{m}^{(t)}$ is learned by maximizing the a posteriori probability given by
$\boldsymbol{\gamma}_{m}^{(t)}=\arg\max_{\boldsymbol{\gamma}_{m}^{(t)}}\quad p\left(\boldsymbol{\gamma}_{m}^{(t)}|\boldsymbol{\mu}_{m}^{(t)}\right),$ (31)
where $p(\boldsymbol{\gamma}_{m}^{(t)}|\boldsymbol{\mu}_{m}^{(t)})$ denotes the marginal posteriori probability of $\boldsymbol{\gamma}_{m}^{(t)}$ given by
$p\left(\boldsymbol{\gamma}_{m}^{(t)}|\boldsymbol{\mu}_{m}^{(t)}\right)=\int p(\mathbf{x}_{m}^{(t)},\boldsymbol{\gamma}_{m}^{(t)},\mathbf{s}_{m}^{(t)}|\boldsymbol{\mu}_{m}^{(t)})\mathrm{d}\mathbf{x}_{m}^{(t)}\mathrm{d}\mathbf{s}_{m}^{(t)}.$ (32)
However, the introduction of the MRF couples the variables $x_{m,n}$, making the integral in (32) computationally intractable. To address this challenge, we propose leveraging GNNs to learn the prior variance parameters. Specifically, using the pairwise MRF model provided in (30), the dependency among the angular-delay channel coefficients $x_{m,n}$ for $\forall m\in\mathcal{M}$ and $\forall n\in\mathcal{N}$ can be represented by an undirected graph $G^{m}=\{V^{m},E^{m}\}$. Here, $V^{m}=\{x_{m,1},\dots,x_{m,n},\dots,x_{m,N}\}$ denotes the set of nodes, and $E^{m}=\{e^{m}_{n,k}\}$, where $n,k\in\mathcal{N}$ and $k\neq n$, represents the set of edges capturing the dependencies. The prior variance $\boldsymbol{\gamma}_{m}$ can then be interpreted as a hidden attribute of $G^{m}$ and effectively learned using GNNs applied to $G^{m}$.
According to the above analysis, the block diagram of the proposed SBL-GNNs is illustrated in Fig. 3, where the network consists of $T$ cascade layers. Each layer shares the same structure, comprising GNNs module and an observation module. The input of the SBL-GNNs network is the received signal of each subarray $\mathbf{y}_{m}$ and the noise level $\zeta_{m}$, with the initial $\boldsymbol{\gamma}_{m}^{(0)}=\mathbf{1}_{N}$ and $\zeta_{m}^{(0)}=1$, and the output is the posterior mean $\boldsymbol{\mu}_{m}^{(T)}$. For any middle layer $t$ with $1<t<T$, its input is the received signal $\mathbf{y}_{m}$, the estimated variance $\boldsymbol{\gamma}_{m}^{(t-1)}$ and noise level $\zeta_{m}^{(t-1)}$ from $t-1$ layer. Similar to the SBL-based algorithms [31, 32], the SBL-GNNs network is sequentially executed until termination by a fixed number of layers.
@@
$\mathbf{a}_{m,n}^{(l)}=\left[\boldsymbol{\lambda}_{m}^{\mathrm{H}}\mathbf{v}_{m,n},v_{m}^{n,n},\alpha_{m},\zeta_{m}\right]^{\mathrm{T}},m\in\mathcal{M},n\in\mathcal{N},$ (33)
@@
$\mathbf{u}^{(0)}_{m,n}=\mathbf{W}_{1}\mathbf{a}_{m,n}+\mathbf{b}_{1},$ (34)
@@ IV-B1 Propagation module
In propagation module, for any variable nodes $x_{m,n}$ and $x_{m,k}$, there exists a factor node, which concentrates the edge attributes and the messages $\mathbf{u}^{(l)}_{m,n}$ and $\mathbf{u}^{(l)}_{m,k}$ from $x_{m,n}$ and $x_{m,k}$ and is given by
$\mathbf{c}_{m}^{n,k}(l)=\mathcal{P}\left(\mathbf{u}^{(l-1)}_{m,n},\mathbf{u}^{(l-1)}_{m,k},\mathbf{f}_{m}^{n,k}\right),$ (35)
@@ IV-B2 Aggregation module
$\mathbf{c}_{m}^{n}(l)=\left[\sum_{n\neq k}\left(\mathbf{c}_{m}^{n,k}(l)\right)^{\mathrm{T}},\mathbf{a}_{m,n}^{\mathrm{T}}\right]^{\mathrm{T}}.$ (36)
@@ IV-B3 Readout module
Since GNNs combine the advantages of DL and MRF in a unique framework to capture the structure of the data into feature vectors that are updated through message-passing among nodes, it has been adopted for MIMO detection task [35, 36, 37] to deal better with the correlated and non-Gaussian residual noise and improving the accuracy of the posterior distribution approximation. However, in existing works, the nodes are usually modeled as users or antennas while the edges among nodes is considered as channels to construct the GNNs. However, the GNNs architecture tailored MIMO detection are inapplicable for the channel estimation since the attributes of graph, e.g., the choices of node and edge, node attributes and edge attributes, need to be reconsidered.
@@ IV-C The Overall Algorithm Description
9:         Compute $\mathbf{u}^{(0)}_{m,n}$ in (34), $n\in\mathcal{N}$;
@@
12:         Compute $\mathbf{c}_{m}^{n,k}(l)$ in (35), $n,k\in\mathcal{N}$ and $k\neq n$;
13:         Compute $\mathbf{c}_{m}^{n}(l)$ in (36), $n\in\mathcal{N}$;
14:         Compute $\mathbf{g}_{m,n}^{(l)}$ in (37), $n\in\mathcal{N}$;
15:         Compute $\mathbf{u}_{m,i}^{(l)}$ in (38), $n\in\mathcal{N}$;
@@
The proposed SBL-GNNs algorithm is composed of two modules, i.e., observation module and GNNs module. The computations in the observation module remain the same as those in the traditional SBL-based algorithm. The difference between SBL-GNNs and traditional SBL is that the update of prior variance is replaced by GNNs. The overall SBL-GNNs algorithm is given in Algorithm. 1. In the following, we elaborate on the computational complexity in terms of the number of multiplications of the proposed SBL-GNNs algorithm. In the observation module, we first need to compute $\boldsymbol{\Phi}_{m}^{\mathrm{H}}\boldsymbol{\Phi}_{m}$ and $\boldsymbol{\Phi}_{m}^{\mathrm{H}}\mathbf{y}_{m}$. The computational complexity is $\mathcal{O}(K^{3}N^{2}PN_{\mathrm{RF}})$ and $\mathcal{O}(K^{2}PN_{\mathrm{RF}}N)$, respectively. Then, the matrix inversion operation in (21) involves the computational complexity of $\mathcal{O}(N^{3}K^{3}+N^{2}K^{2}+2NK)$. In the GNNs module, the main computational complexity is from the MLP, whose complexity is $\mathcal{O}(LN_{\mathrm{u}}N_{\mathrm{h}_{1}}N_{\mathrm{h}_{2}})$. Thus, the over computational complexity is $\mathcal{O}(K^{3}N^{2}PN_{\mathrm{RF}}T+K^{2}PN_{\mathrm{RF}}NT+N^{3}K^{3}T+N^{2}K^{2}T+2NKT+LN_{\mathrm{u}}N_{\mathrm{h}_{1}}N_{\mathrm{h}_{2}}T)$.
@@ IV-D Training Scheme
The proposed SBLNet estimator was implemented in Python using PyTorch. The training process consisted of 50 epochs, with each epoch comprising 320 batches of 64 samples. Each sample included realizations of $\mathbf{x}_{m}$, $\boldsymbol{\Phi}_{m}$, $\mathbf{y}m$, and $\mathbf{n}m$ generated according to (11). The signal-to-noise ratio (SNR) values were uniformly distributed in the range $[SNR{\mathrm{min}},SNR{\mathrm{max}}]$. The Adam optimizer was employed for training with a learning rate of $1\times 10^{-4}$. A mean squared error (MSE) loss function was utilized to update the learnable parameters via the back-propagation algorithm.
@@
V Central Bayesian Refinement Algorithm
@@
$=\prod_{j=1}^{N_{t}}p(x_{j}|\upsilon_{j})p(\upsilon_{j}|o_{j})p(\mathbf{o}),$
@@
The Markov chain-based prior model facilitates a flexible and effective representation of sparsity and correlations among the coefficients $x_{j}$. Specifically, the precision parameters of the variance, $\upsilon_{j}$, regulate the sparsity of $x_{j}$, where $\upsilon_{j}^{-1}$ approaching zero effectively enforces $x_{j}$ to be zero. Furthermore, the Markov chain models the variance states, encouraging structured block patterns while suppressing isolated coefficients that deviate from their neighboring values. In particular, the state of each coefficient being non-zero is conditioned on the states of its adjacent coefficients, with transition probabilities $p(o_{j}|o_{j-1})$ dictating the likelihood of transitions between states. When a coefficient is in the non-zero state, the transition probability of remaining in this state is expected to be higher, thereby promoting continuity and reducing the occurrence of isolated non-zero coefficients.
@@
The dependencies of the random variables in the factorization (42) can be shown by a factor graph as depicted in Fig. 4, where $f_{r_{j}}\triangleq p(r_{j}\mid x_{j},\kappa)$, $f_{x_{j}}\triangleq p(x_{j}\mid\upsilon_{j})$, $f_{\upsilon_{j}}\triangleq p(\upsilon_{j}\mid o_{j})$, and $f_{\kappa}\triangleq p(\kappa)$. To achieve a computationally efficient inference of problem (42), we propose a variational message passing based method. In the following, we derive the forward (from left to right) and backward (from right to left) message passing in detail.
@@ V-2 Backward Message Passing
The refinement scheme in [24] relies solely on angular-delay sparsity, ignoring coefficient correlations. It identifies significant coefficients and suppresses others using soft-thresholding window. Additionally, the method depends heavily on precise covariance matrix knowledge, limiting its practical applicability. In contrast, this work introduces a Bayesian denoising approach that jointly exploits angular-delay sparsity and coefficient correlations via a Markov chain-based hierarchical prior, significantly enhancing refinement performance.
@@ VI Simulation Results
In this section, we evaluate the performance of the proposed SBL-GNNs and channel refinement algorithms. In the simulation, the key parameters are provided in Table I. In particular, we consider normalized mean square error (NMSE) as performance metrics, which is defined as $\mathrm{NMSE}\triangleq{\lVert\hat{\mathbf{H}}_{m}(\tilde{\mathbf{H}})-\mathbf{H}_{m}(\mathbf{H})\rVert^{2}_{\mathrm{F}}}/{\lVert\mathbf{H}_{m}(\mathbf{H})\lVert^{2}_{\mathrm{F}}}$, where $\mathbf{H}_{m}(\mathbf{H})$ and $\hat{\mathbf{H}}_{m}(\tilde{\mathbf{H}})$ are the true channel and estimated channel, respectively. In addition, the received SNR is defined as $10\log_{10}\left(\lVert\mathbf{WH}\rVert^{2}_{\mathrm{F}}/\lVert\mathbf{N}\rVert^{2}_{\mathrm{F}}\right)$. To evaluate channel estimation performance in LPUs, we compare our proposed SBL-GNNs algorithm with the following benchmarks, which adopt different sparse prior models:
  • UAMP-SBL [13]: Unity approximate message passing algorithm with a 4-connect MRF prior to characterize the dependencies among neighboring coefficients.
@@
Subsequently, we first investigate the effects of the number of subarrays on estimation performance and evaluate the algorithm computational complexity through algorithm runtime. Then, we evaluate the effectiveness of the SBL-GNNs and channel refinement algorithms, respectively.

### v4

@@ IV-B Proposed SBL-GNNs Estimator
Utilizing the Lemma 1, the precision vector $\boldsymbol{\gamma}_{m}^{(t)}$ is learned by maximizing a posteriori probability given by
$\boldsymbol{\gamma}_{m}^{(t+1)}=\arg\max_{\boldsymbol{\gamma}_{m}}\quad p\left(\boldsymbol{\gamma}_{m}|\boldsymbol{\mu}_{m}^{(t)}\right),m\in\mathcal{M},$ (30)
where $p(\boldsymbol{\gamma}_{m}|\boldsymbol{\mu}_{m}^{(t)})$ denotes the marginal posteriori probability of $\boldsymbol{\gamma}_{m}$ given by
$p\left(\boldsymbol{\gamma}_{m}|\boldsymbol{\mu}_{m}^{(t)}\right)=\int p(\mathbf{x}_{m},\boldsymbol{\gamma}_{m},\mathbf{s}_{m}|\boldsymbol{\mu}_{m}^{(t)})\mathrm{d}\mathbf{x}_{m}\mathrm{d}\mathbf{s}_{m}.$ (31)
However, the introduction of the MRF couples the variables $x_{m,n}$, making the integral in (31) computationally intractable. To address this challenge, we propose leveraging GNNs to learn the prior variance parameters. Specifically, using the pairwise MRF model provided in (29), the dependency among the angular-delay channel coefficients $x_{m,n}$ for $\forall m\in\mathcal{M}$ and $\forall n\in\mathcal{V}$ can be represented by an undirected graph $G^{m}=\{V^{m},E^{m}\}$. Here, $V^{m}=\{x_{m,1},x_{m,2},\dots,x_{m,NK}\}$ denotes the set of nodes, and $E^{m}=\{e^{m}_{n,k}\}$, where $n,k\in\mathcal{V}$ and $k\neq n$, represents the set of edges capturing the dependencies. The prior variance $\boldsymbol{\gamma}_{m}$ can then be interpreted as a hidden attribute of $G^{m}$ and effectively learned using GNNs applied to $G^{m}$.
According to the above analysis, the block diagram of the proposed SBL-GNNs is illustrated in Fig. 3, where the network consists of $T$ cascade layers. Each layer shares the same structure, comprising GNNs module and an observation module. The input of the SBL-GNNs network is the received signal of each subarray $\mathbf{y}_{m}$ and the noise level $\zeta_{m}$, with the initial $\boldsymbol{\gamma}_{m}^{(0)}=\mathbf{1}_{NK}$ and $\zeta_{m}^{(0)}=1$, and the output is the posterior mean $\boldsymbol{\mu}_{m}^{(T)}$. For any middle layer $t$ with $1<t<T$, its input is the received signal $\mathbf{y}_{m}$, the estimated variance $\boldsymbol{\gamma}_{m}^{(t-1)}$ and noise level $\zeta_{m}^{(t-1)}$ from $t-1$ layer. Similar to the SBL-based algorithms [34, 35], the SBL-GNNs network is sequentially executed until termination by a fixed number of layers.
@@
$\mathbf{a}_{m,n}^{(l)}=\left[\boldsymbol{\lambda}_{m}^{\mathrm{H}}\mathbf{v}_{m,n},v_{m}^{n,n},\alpha_{m},\zeta_{m}\right]^{\mathrm{T}},m\in\mathcal{M},n\in\mathcal{V},$ (32)
@@
$\mathbf{u}^{(0)}_{m,n}=\mathbf{W}_{1}\mathbf{a}_{m,n}+\mathbf{b}_{1},m\in\mathcal{M},n\in\mathcal{V},$ (33)
@@ IV-B1 Propagation module
In propagation module, for any variable nodes $x_{m,n}$ and $x_{m,k}$ with $n\neq k$, there exists a factor node, which concentrates the edge attributes and the messages $\mathbf{u}^{(l)}_{m,n}$ and $\mathbf{u}^{(l)}_{m,k}$ from $x_{m,n}$ and $x_{m,k}$ and is given by
$\mathbf{c}_{m}^{n,k}(l)=\mathcal{P}\left(\mathbf{u}^{(l-1)}_{m,n},\mathbf{u}^{(l-1)}_{m,k},\mathbf{f}_{m}^{n,k}\right),m\in\mathcal{M},n,k\in\mathcal{V},$ (34)
@@ IV-B2 Aggregation module
$\mathbf{c}_{m}^{n}(l)=\left[\sum_{k\neq n}\left(\mathbf{c}_{m}^{n,k}(l)\right)^{\mathrm{T}},\mathbf{a}_{m,n}^{\mathrm{T}}\right]^{\mathrm{T}},$ (35)
@@ IV-B3 Readout module
Due to the differences in estimation and prior distribution, the pair-wise MRF representation exhibits significant differences with existing works in [38, 39, 40], resulting in distinct graph attributes and GNN architectures. Thus, we reconsider the graph attributes, including the pair-wise MRF representation (see (29)) as well as the node and edge attributes tailored to the channel estimation problem. Meanwhile, from a system perspective, the proposed GNN is designed to capture the correlations across different channel coefficients through the fully connected MRF, whereas the GNNs in [38, 39, 40] primarily address correlated and non-Gaussian residual noise. Additionally, due to the differences in the types of parameters to be estimated, the readout module of GNNs differ fundamentally from existing works. Specifically, the readout module designed for discrete variables in MIMO detection tasks is not applicable to channel estimation. To this end, we redesigned the readout module to accommodate the regression tasks associated with continuous variables in the channel estimation problem.
@@ IV-C The Overall Algorithm Description
9:         Compute $\mathbf{u}^{(0)}_{m,n}$ in (33), $n\in\mathcal{V}$;
@@
12:         Compute $\mathbf{c}_{m}^{n,k}(l)$ in (34), $n,k\in\mathcal{V}$ and $k\neq n$;
13:         Compute $\mathbf{c}_{m}^{n}(l)$ in (35), $n\in\mathcal{V}$;
14:         Compute $\mathbf{g}_{m,n}^{(l)}$ in (36), $n\in\mathcal{V}$;
15:         Compute $\mathbf{u}_{m,i}^{(l)}$ in (37), $n\in\mathcal{V}$;
@@
The proposed SBL-GNNs algorithm is composed of two modules, i.e., observation module and GNNs module. The computations in the observation module remain the same as those in the traditional SBL-based algorithm. The difference between SBL-GNNs and traditional SBL is that the update of prior variance is replaced by GNNs. The overall SBL-GNNs algorithm is given in Algorithm 1. In the following, we elaborate on the computational complexity in terms of the number of multiplications of the proposed SBL-GNNs algorithm. In the observation module, we first need to compute $\boldsymbol{\Phi}_{m}^{\mathrm{H}}\boldsymbol{\Phi}_{m}$ and $\boldsymbol{\Phi}_{m}^{\mathrm{H}}\mathbf{y}_{m}$. The computational complexity is $\mathcal{O}(K^{3}N^{2}PN_{\mathrm{RF}})$ and $\mathcal{O}(K^{2}PN_{\mathrm{RF}}N)$, respectively. Then, the matrix inversion operation in (21) involves the computational complexity of $\mathcal{O}(N^{3}K^{3}+N^{2}K^{2}+2NK)$. In the GNNs module, the main computational complexity is from the MLP, whose complexity is $\mathcal{O}(LN_{\mathrm{u}}N_{\mathrm{h}_{1}}N_{\mathrm{h}_{2}})$. Thus, the over computational complexity is $\mathcal{O}(K^{3}N^{2}PN_{\mathrm{RF}}T+K^{2}PN_{\mathrm{RF}}NT+N^{3}K^{3}T+N^{2}K^{2}T+2NKT+LN_{\mathrm{u}}N_{\mathrm{h}_{1}}N_{\mathrm{h}_{2}}T)$.
@@ IV-D Training Scheme
The proposed SBL-GNNs estimator was implemented in Python using PyTorch. The training process consisted of 50 epochs, with each epoch comprising 320 batches of 64 samples. Each sample included realizations of $\mathbf{x}_{m}$, $\boldsymbol{\Phi}_{m}$, $\mathbf{y}m$, and $\mathbf{n}m$ generated according to (11). The signal-to-noise ratio (SNR) values were uniformly distributed in the range $[SNR{\mathrm{min}},SNR{\mathrm{max}}]$. The Adam optimizer was employed for training with a learning rate of $1\times 10^{-4}$. A mean squared error (MSE) loss function was utilized to update the learnable parameters via the back-propagation algorithm.
@@
V Proposed Centralized Bayesian Refinement Algorithm
@@
$=\prod_{j\in\mathcal{J}}p(x_{j}|\upsilon_{j})p(\upsilon_{j}|o_{j})p(\mathbf{o}),$
@@
According to (39) and (40), the prior of $x_{j}$ is given by
$p(x_{j})=\int\mathcal{CN}(x_{j};0,\upsilon_{j}^{-1})p(\upsilon_{j}|o_{j})p(\mathbf{o})\mathrm{d}\upsilon_{j}\mathrm{d}\mathbf{o},j\in\mathcal{J}.$ (41)
From (41), the Markov chain-based prior model facilitates a flexible and effective representation of sparsity and correlations among the coefficients $x_{j}$. Specifically, the precision parameters of the variance, $\upsilon_{j}$, regulate the sparsity of $x_{j}$, where $\upsilon_{j}$ approaching zero effectively enforces $x_{j}$ to be zero. Furthermore, the Markov chain models the variance states, encouraging structured block patterns while suppressing isolated coefficients that deviate from their neighboring values. Precisely, the state of each coefficient being non-zero is conditioned on the states of its adjacent coefficients, with transition probabilities $p(o_{j}|o_{j-1})$ dictating the likelihood of transitions between states. When a coefficient is in the non-zero state, the transition probability of remaining in this state is expected to be higher, thereby promoting continuity and reducing the occurrence of isolated non-zero coefficients.
@@
The dependencies of the random variables in the factorization (42) can be shown by a factor graph as depicted in Fig. 4, where $f_{r_{j}}\triangleq p(r_{j}\mid x_{j},\kappa)$, $f_{x_{j}}\triangleq p(x_{j}\mid\upsilon_{j})$, $f_{\upsilon_{j}}\triangleq p(\upsilon_{j}\mid o_{j})$, and $f_{\kappa}\triangleq p(\kappa)$. To achieve a computationally efficient inference of problem (42), we propose a variational message passing and belief propagation based refinement algorithm.
Remark 4.
From (41), it is evident that, unlike [37], where each $x_{j}$ is assigned an independent variance precision parameter, our approach explicitly captures the correlations among angular-delay channel coefficients by the Markov chain. This fundamental distinction leads to significant differences in the factor graph structure. Specifically, the sparse prior models in [37] are typically formulated for loop-free factor graphs. In contrast, our approach introduces loops and higher-order dependencies through the Markov chain, resulting in a more intricate factor graph representation, as illustrated in Fig. 4. This structural complexity necessitates distinct message update equations at the variable nodes, deviating from conventional message-passing strategies designed for [37].
In the following, we derive the forward (from left to right) and backward (from right to left) message passing in detail.
@@ V-2 Backward Message Passing
The refinement scheme in [30] relies solely on angular-delay sparsity, ignoring coefficient correlations. It identifies significant coefficients and suppresses others using soft-threshold window. Additionally, the method depends heavily on precise covariance matrix knowledge, limiting its practical applicability. In contrast, this work introduces a Bayesian denoising approach that jointly exploits angular-delay sparsity and coefficient correlations via a Markov chain-based hierarchical prior, enhancing refinement performance.
@@ VI Simulation Results
In this section, we evaluate the performance of the proposed SBL-GNNs and channel refinement algorithms. In the simulation, the key parameters are provided in Table I. In particular, we consider NMSE as performance metrics, which is defined as $\mathrm{NMSE}\triangleq{\lVert\hat{\mathbf{H}}_{m}(\tilde{\mathbf{H}})-\mathbf{H}_{m}(\mathbf{H})\rVert^{2}_{\mathrm{F}}}/{\lVert\mathbf{H}_{m}(\mathbf{H})\lVert^{2}_{\mathrm{F}}}$, where $\mathbf{H}_{m}(\mathbf{H})$ and $\hat{\mathbf{H}}_{m}(\tilde{\mathbf{H}})$ are the true channel and estimated channel, respectively. In addition, the received SNR is defined as $10\log_{10}\left(\lVert\mathbf{WH}\rVert^{2}_{\mathrm{F}}/\lVert\mathbf{N}\rVert^{2}_{\mathrm{F}}\right)$. To evaluate channel estimation performance in LPUs, we compare our proposed SBL-GNNs algorithm with the following benchmarks, which adopt different sparse prior models:
  • UAMP-MRF [17]: Unity approximate message passing algorithm with a 4-connect MRF prior to characterize the dependencies among neighboring coefficients.
TABLE II: Computational Complexity Comparison
Algorithms Computational complexity $\mathcal{O}(\cdot)$
StdSBL $\mathcal{O}(N^{3}K^{3}T_{1})$
PC-SBL $\mathcal{O}(N^{3}K^{3}T_{2})$
VSP $\mathcal{O}(N^{3}K^{3}T_{3})$
UAMP $\mathcal{O}(K^{3}N^{2}PN_{\mathrm{RF}}T_{4})$
Proposed SBL-GNNs
$\mathcal{O}(K^{3}N^{2}PN_{\mathrm{RF}}T+K^{2}PN_{\mathrm{RF}}NT$
$+N^{3}K^{3}T+LN_{\mathrm{u}}N_{\mathrm{h}_{1}}N_{\mathrm{h}_{2}}T)$
The computational complexities are detailed in Table II, where $T_{1}$, $T_{2}$, $T_{3}$, $T_{4}$, and $T$ denote the number of iterations required for different algorithms, respectively. From Table II, it can be observed that, for each iteration, the StdSBL, PC-SBL, VSP, and the proposed SBL-GNNs algorithms have comparable computational complexities due to the involvement of matrix inversion operations. In contrast, the UAMP algorithm, which only involves matrix product operations, exhibits the lowest computational complexity per iteration. Notably, owing to the powerful capability of GNNs in capturing the correlation and sparsity of the channel, the SBL-GNNs algorithm requires significantly fewer iterations to achieve satisfactory performance, i.e., $T\ll T_{1},T_{2},T_{3},T_{4}$. Consequently, the proposed SBL-GNNs algorithm exhibits a lower overall computational complexity, as will be validated later.
Figure 5: NMSE of SBL-GNNs versus subarray number.